OpenStudy (anonymous):

(2^2)(3^3)(4^4)(5^5)(6^6)(7^7)(8^8)(9^9)(10^10)=? ?=

OpenStudy (jamesj):

A large number. A more interesting question might be: without trying to calculate the number explicitly, how many zeros does it have at the end?

OpenStudy (anonymous):

???

OpenStudy (jamesj):

215779412229418562091680268288000000000000000

OpenStudy (anonymous):

o0 0

OpenStudy (anonymous):

I think i am mind struck

OpenStudy (jamesj):

BTW, the answer to the question, how many zeros, as you see if 15. But you can calculate it this way: how many zeros doest the integer p have at the end? # of zeros = # of factors of 10 = # of factors of 2 with matching # of factors of 5 In your number there are 5 factors of 5 from 5^5, and then another 10 factors of 5 from 10^10. Hence in total 15 factors of 5. there are certainly at least that many factors of 2, as they come from 2^2, 4^4, 6^6, 8^8 and 10^10. Hence 15 = 10 + 5 = # of factors of 2 with matching # of factors of 5 = # of factors of 10 = # of zeros

OpenStudy (anonymous):

THX, i learned alot!